Project

General

Profile

NDOS Commissioning Meeting 2011/04/04

Attendance: Jon, Chad, Mark Messier, Peter, Andrew, Mat, Leon, Alena, Tom C, Kurt, Gavin, Minerba, Tricia, Art, TJ

Agenda:

  1. NDOS Status / Installation Plans
  2. DAQ/Electronics Status & Plans
  3. DCS Status & Plans
  4. AOB

NDOS Status / Installation Plans:

  • Mat:
    • Karen is continuing with the filling operation.
    • The vent-port filling head didn't work as planned.
    • Were able to get modified nozzle head to fill some modules that couldn't be reached..
    • Will finish off verticals by tomorrow afternoon. Still no procedure for topping off horizontals.
    • Hooked five new DCMs into timing chain. Master-slave TDU has not been tested.
  • Should DCM timing tests be part of DCM checkout? Leon: yes. But this is not part of Sue's DCM validation script.
  • Andrew: hardware supports ability to verify timing data, but software doesn't exist yet.
  • Andrew: spoke to Rick and Vince re: checkout of TDU master-slave operation. They'd like Ron R. to work on this. Peter and Leon
  • Mat: do not have data cables run for muon catcher because there is no map for how we'll hook up FEBs on side DCM. Andrew will work on that.
  • Mat: would like FEB validation script run on DCM 3-3-1, would like to install APD tomorrow. Peter: can be done today.
  • Leon: still on schedule for CAP boards.
  • Mark: still missing ~26 FEBs. Mat: lower priority. Cable-ectomy is probably best approach, but list needs to be reviewed.

DAQ/Electronics Status & Plans:

  • Peter: need to update development software today
  • Gennadiy found that missing triggers are not due to DDS.
  • Qiming started work on watchdog program
  • Peter created an ntuple-maker that mines Ganglia data and DCS data.

DCS Status & Plans:

  • Athans: testing apps on test-stand since last week. Found DDS server has been crashing every ~12 hours. Will try
  • Chad: will install RTDs this week. Will begin this afternoon; patch panels first, will connect existing devices to panel (will need to rewire some).

AOB

  • daqlogs disk was full on Friday; DAQ group will work on a cron-job to take care of this.